OpenStudy (anonymous):

SOMEONE PLLLZZ HELP!!!!!!!! Solve the equation. Choose the correct solution. (x-5/x)-2=(x+3/x)

OpenStudy (anonymous):

\[\frac{x-5}{x}-2=\frac{x+3}{x}\]?

OpenStudy (anonymous):

i mean is that the problem?

OpenStudy (anonymous):

yes that is it

OpenStudy (anonymous):

ok good first step is to subtract on the left

OpenStudy (anonymous):

\[\frac{x-5}{x}-2=\frac{x-5-2x}{x}=\frac{-x-5}{x}\]

OpenStudy (anonymous):

then you get \[\frac{-x-5}{x}=\frac{x+3}{x}\] and now that the denominators are the same, all that is left to solve is \[-x-5=x+3\]

OpenStudy (anonymous):

ok so the answer would be 8

OpenStudy (anonymous):

no it would be -4 thnx for the help